# Scalable Approximate Biclique Counting over Large Bipartite Graphs

### v1

@@ 8       Initialize $\kappa\leftarrow 1$;
10             Initialize an array $Cnt$ with zeros;
@@ 13                   $c(u)\leftarrow\kappa$;
14                   Update $Cnt$;
15                   if $\exists x,Cnt[x]\geq q$ then
16                         Revert $Cnt$’s change;
@@ 4.2. Pattern Counting
After coloring, we enhance the motif we are counting with a condition of distinct color. However, it is still hard to count bicliques directly. Instead, we count the number of $(p,q)$-brooms. Since we assign a concrete integer for each color in Algorithm 1 for any $(p,q)$-biclique, to avoid overcounting, we assume its vertex ordering of $U^{\prime}$ and $V^{\prime}$ is following the color from small to large. With this exact ordering, by Definition 4.1, there is exactly one $(p,q)$-broom subgraph in each $(p,q)$-biclique. Note that the color of these $(p,q)$-brooms is also unique, which is convenient for counting. Algorithm 2 is able to compute the number of $(p,q)$-brooms of this type. We then show how to use this value to approximate the biclique count.
The whole process is dynamic programming where its state is represented by the total number of edges we have accumulated so far and the last edge we keep track of. To well define this dynamic programming state, we need to specify a few more properties in this type of $(p,q)$-broom:
A. Any $(p,q)$-broom $H$ is connected and contains exactly $p+q-1$ edges. That is to say, it is actually a tree.
@@
For (2) and (3), we signal them with a ”direction” as either Up or Down. For example, as shown in the $(6,3)$-broom in Figure 3, $(u_{1},v_{1})$ is with the smallest partial order. Its direction is Down since it shares the same endpoint $v_{1}$ with the next edge $(u_{2},v_{1})$. Similarly, $(u_{2},v_{1})$, $(u_{3},v_{2})$, $(u_{4},v_{2})$, $(u_{5},v_{3})$ are also with direction Down, highlighted with blue arrows. $(u_{3},v_{1})$, $(u_{5},v_{2})$ are with direction Up, highlighted with reversed red arrows. $(u_{6},v_{3})$ has no direction since it is the edge with the maximum tuple.
@@
[…]
@@
Output: $Dp$: a $2D$ array; $B$: the total number of $(p,q)$-brooms in $G$.
@@ 3 Sort $E(G)$ by the increasing order of $(c(u),c(v))$;
5Initialize $Dp$ with zeros;
@@ 4.3. Approximate Counting via Sampling
After CountingIndex$(G,p,q)$, we have accumulated $B$ $(p,q)$-brooms. To approximate the quantity relation between our computed $(p,q)$-brooms and $(p,q)$-bicliques, we can do the following sampling process:
@@
After applying this sampling process sufficiently many times, we can roughly approximate the number of $(p,q)$-bicliques by $cnt_{biclique}/cnt_{broom}\times B$.
However, such a sampling process is too inefficient. We now propose Algorithm 3 to accelerate this process, utilizing the computed dynamic programming results from CountingIndex$(G,p,q)$. Recall that $Dp[t][(u,v)]$ represents the number of ways to build a growing $(p,q)$-broom til the $t$-th edge, which is $(u,v)$.
@@
To begin with, we start by sampling the last edge of the $(p,q)$-broom, denoted as $(u_{last},v_{last})$ (Algorithm 3). Here, the sampling should follow the weight distribution of $\{Dp[p+q-1][\cdot]\}$, indicating how many $(p,q)$-brooms end with each edge. The following process is similar to reverting the dynamic programming process. Instead of building the $(p,q)$-broom from the $1$-st edge to the $(p+q-1)$-th edge, we do it reversely. We use $(u_{cur},v_{cur})$, $U^{\prime}$, $V^{\prime}$ to denote the current growing $(p,q)$-broom. They are initialized as $(u_{last},v_{last})$, $\{u_{last}\}$, $\{v_{last}\}$ respectively (Algorithm 3, 3). Specifically, since $(u_{last},v_{last})$ is the $(p+q-1)$-th edge, we are now adding the $(p+q-2)$-th edge til the $1$-th edge gradually (Algorithm 3). Assume we are processing the $i$-th edge now, and we know the $(i+1)$-th edge is $(u_{cur},v_{cur})$. We initialize an edge set $S$ to store the candidate edge for the $i$-th edge. Based on the direction signaling, we can know whether it should share the common endpoint in $U$ or $V$. If $Dir[i]$ is Up (Algorithm 3), then we should find all $w\in N(u_{cur},G)$ such that $c(w)<c(v_{cur})$ and $U^{\prime}\subseteq N(w,G)$. We assign $S$ to be $\{(u_{cur},w)\}$ (Algorithm 3). The first condition is for obeying the color order, and the second condition indicates that adding $w$ can still guarantee that the induced subgraph is biclique. Similarly, if $Dir[i]$ is Down (Algorithm 3), we find all all $w\in N(v_{cur},G)$ such that $c(w)<c(u_{cur})$ and $V^{\prime}\subseteq N(w,G)$. Then we assign $S$ to be $\{(w,v_{cur})\}$ (Algorithm 3).
@@
The idea to accelerate the sampling is, by building the $(p,q)$-broom, we strictly guarantee that it will correspond to a $(p,q)$-biclique while keeping track of the probability of sampling out this $(p,q)$-broom, which can be computed through the dynamic programming table. In the beginning, we initialize $ans$ to be $1$ (Algorithm 3). Then, for the $i$-th edge, the probability contribution of sampling it out from $S$ should be the number of the growing broom ends with $(u,v)\in S$ divided by the total number of possible brooms at this step, which is $\frac{\sum_{(u,v)\in S}Dp[i][(u,v)]}{Dp[i+1][(u_{cur},v_{cur})]}$. We multiply $ans$ by this value (Algorithm 3). If $ans$ becomes 00, then there is no possible $(p,q)$-biclique from the current sampled $(p,q)$-broom. We return with 00 in this case (Algorithm 3). Now we are ready to sample the $i$-th edge from $S$ (Algorithm 3). We do so following the normalized distribution of possible brooms from this step ($\{Dp[i][\cdot]\}$). To continue the process, we assign $(u_{cur},v_{cur})$, $U^{\prime}$, $V^{\prime}$ to $(u_{next},v_{next})$, $U^{\prime}\cup\{u_{next}\}$, $V^{\prime}\cup\{v_{next}\}$, respectively (Algorithm 3, 3).
In the end, we successfully sample a $(p,q)$-broom that corresponds to one $(p,q)$-biclique and compute the probability of sampling it out. We return this probability by multiplying $B$ as the approximate count of $(p,q)$-bicliques (Algorithm 3).
@@
It is not hard to see that the time complexity is dominated by the cost of updating $S$, which requires enumerating all neighbors. The loop only lasts $O(p+q)$. Therefore, the time complexity is $O((p+q)\times\Delta)$, where $\Delta$ denotes the maximal degree in $G$.
@@
Input: $G$: a colored bipartite graph; $p,q$: two parameters; $Dp$: a $2D$ array; $B$: the total number of $(p,q)$-brooms in $G$.
Output: $ans$: estimate number of $(p,q)$-clique in $G$.
@@ 19       $U^{\prime}\leftarrow U^{\prime}\cup\{u_{next}\},V^{\prime}\leftarrow V^{\prime}\cup\{v_{next}\}$;
Overall Algorithm. We provide Algorithm 4 as a black box to call all three subroutines properly and output the estimated answer of the biclique count. We start by coloring (Algorithm 4) and pattern counting (Algorithm 4). Then by the input sampling size parameter $T$, we repeatedly call Sampling$(G,p,q,Dp,B)$ and aggregate the return to $\hat{C}$. The output approximate $(p,q)$-biclique count is the average of all $T$ attempts (Algorithm 4). Note that as a common trick, we will execute core-reduction for the original graph (Yang et al., 2021): For any query $(p,q)$, we split the graph and reduce the query pair to $(p-1,q)$.
@@
1 $G\leftarrow$ Coloring$(G,p,q)$;
2 $Dp,B\leftarrow$ CountingIndex$(G,p,q)$;
@@ 5       $\hat{C}\leftarrow\hat{C}\ +$ Sampling$(G,p,q,Dp,B)$;
Unbiasedness. We now show that the CBS method is unbiased. Let $(u_{cur},v_{cur})$, $U^{\prime}$, and $V^{\prime}$ represent the current growing $(p,q)$-broom. For clarity, we introduce the following definitions:
  • $\mathcal{F}_{U^{\prime},V^{\prime}}$: The total value to be multiplied into $ans$, defined as the product of the last $i=(p+q)-(|U^{\prime}|+|V^{\prime}|)$ fractions.
@@
(a) When $|U^{\prime}|+|V^{\prime}|=p+q$, we have $\mathcal{B}_{U^{\prime},V^{\prime}}=1$. In this case, $\mathcal{F}_{U^{\prime},V^{\prime}}=\mathcal{C}_{U^{\prime},V^{\prime}}=1$ if $U^{\prime},V^{\prime}$ forms a $(p,q)$-biclique; otherwise $\mathcal{F}_{U^{\prime},V^{\prime}}=\mathcal{C}_{U^{\prime},V^{\prime}}=0$. Hence, $\mathbb{E}\left[\mathcal{F}_{U^{\prime},V^{\prime}}\right]=\mathcal{C}_{U^{\prime},V^{\prime}}/\mathcal{B}_{U^{\prime},V^{\prime}}$ trivially holds in this case. (b) Suppose that the lemma holds for all growing $(p,q)$-brooms with $\ |U^{\prime}|+|V^{\prime}|=k+1$ For any broom with $|U^{\prime}|+|V^{\prime}|=k$, let $P=\frac{\sum_{(u,v)\in S}Dp[i][(u,v)]}{Dp[i+1][(u_{cur},v_{cur})]}$. We then have
@@
Since $u_{cur}=u$ or $v_{cur}=v$ always holds, we have that $|U^{\prime}\cup\{u\}|+|V^{\prime}\cup\{v\}|=|U|+|V|+1=k+1$. According to the definition of $Dp$ and $\mathcal{B}$, We can derive that $Dp[i][(u,v)]=\mathcal{B}_{U^{\prime}\cup\{u\},V^{\prime}\cup\{v\}}$, and thus
@@
$=\frac{\sum\limits_{(u,v)\in S}\mathcal{C}_{U^{\prime}\cup\{u\},V^{\prime}\cup\{v\}}}{\mathcal{B}_{U^{\prime},V^{\prime}}}$
$=\frac{\mathcal{C}_{U^{\prime},V^{\prime}}}{\mathcal{B}_{U^{\prime},V^{\prime}}},$
@@
Let $ans_{i}$ be the value of $ans$ in the $i$-th sampling. Let $\hat{C}=\frac{1}{T}\sum_{i=1}^{T}ans_{i}\times B$. Then $\hat{C}$ is an unbiased estimator of the number of $(p,q)$-bicliques in $G$.
@@
Then we can derive that: $\mathbb{E}\left[\hat{C}\right]=\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}\left[ans_{i}\times B\right]=\mathcal{C}_{\emptyset,\emptyset}$. Note that $\mathcal{C}_{\emptyset,\emptyset}$ is the number of $(p,q)$-bicliques in $G$ according to the definition. Therefore, $\hat{C}$ is an unbiased estimator of the number of $(p,q)$-bicliques in $G$. ∎
@@
Error Analysis. we now analyze the estimation error of our sampling algorithms. Our analysis relies on the classic Hoeffding’s inequalities, which are shown below.
@@
For the random variables $X_{i}\in[0,M]$,$1\leq i\leq n$, we let $X=\sum_{i=1}^{n}X_{i}$. Then for $\epsilon>0$, we have
@@
We can derive $ans_{i}\times B\in[0,B]$ since the values multiplied to $ans$ are all probabilities in $[0,1]$. And we have $\mathbb{E}\left[\hat{C}T\right]=CT$ based on Theorem 4.4. Given a positive value $\epsilon$, applying Lemma 4.5 by plugging in $n=T$, $X_{i}=ans_{i}\times B$, $X=\sum_{i=1}^{n}X_{i}=\hat{C}T$, we then have
@@
Further, we have
$\Pr(\frac{|\hat{C}-C|}{C}\geq\epsilon)\leq 2\exp(-\frac{2\epsilon^{2}C^{2}T}{B^{2}})$
Let $2\exp(-\frac{2\epsilon^{2}C^{2}T}{B^{2}})\leq\alpha$, we can derive that $T\geq\frac{B^{2}}{2\epsilon^{2}C^{2}}\ln(\frac{2}{\alpha})$. ∎
From Theorem 4.6, the sample size $T$ is mainly determined by $(\frac{B}{C})^{2}$. That is to say, we need a larger sample size $T$ to ensure high accuracy with a larger $(\frac{B}{C})^{2}$, where $B$ is the number of $(p,q)$-brooms, $C$ is the number of $(p,q)$-bicliques. As shown in Section A.3, $(\frac{B}{C})^{2}$ is usually small. Therefore, our algorithm generally does not need a large $T$ to achieve good accuracy in real-world datasets.

### v2

@@ 8  Initialize $\kappa\leftarrow 1$;
10     Initialize an array CntCnt with zeros;
@@ 13        $c(u)\leftarrow\kappa$;
14        Update CntCnt;
15        if $\exists x,Cnt[x]\geq q$ and $c(x)=\kappa$ then
16           Revert CntCnt’s change;
@@ 21     $S\leftarrow S_{next}$;
Randomization is a classical way to produce a solution in graph coloring. We have conducted an additional experiment in Section 6.1 using three different ordering strategies: ascending order, descending order, and random order. The results indicate that there are no significant differences in efficiency or error between the three strategies. There is also another ablation study on the effectiveness of adapting the coloring process in counting bicliques. As shown in Section 6.1, with a slight increase in runtime, the coloring process enhances the accuracy in turn.
@@ 4.2. Pattern Counting
After coloring, we enhance the motif we are counting with a condition of distinct colors. However, it is still hard to count bicliques directly. Instead, we count the number of $(p,q)$-brooms. Since we assign a concrete integer for each color in Algorithm 1 for any $(p,q)$-biclique, to avoid overcounting, we assume its vertex ordering of $U^{\prime}$ and $V^{\prime}$ is following the color from small to large. With this exact ordering, by Definition 4.1, there is exactly one $(p,q)$-broom subgraph in each $(p,q)$-biclique. Note that the color of these $(p,q)$-brooms is also unique, which is convenient for counting. Algorithm 2 is able to compute the number of $(p,q)$-brooms of this type. We then show how to use this value to approximate the biclique count.
The whole process is a dynamic programming approach, where its state is represented by the total number of edges accumulated so far and the last edge being tracked. To well define this dynamic programming state, we need to specify a few more properties in this type of $(p,q)$-broom:
A. Any $(p,q)$-broom $H$ is connected and contains exactly p+q1p+q-1 edges. That is to say, it is actually a tree.
@@
Here, for sorting, the comparison between two tuples $(c(u),c(v))$ and $(c(u^{\prime}),c(v^{\prime}))$ is first done by verifying if $c(u)<c(u^{\prime})).$ If $c(u)=c(u^{\prime})$, we then check if $c(v)<c(v^{\prime})$. Therefore, the maximum tuple refers to the last element after sorting in this manner. If For (2) and (3), we signal them with a ”direction” as either Up or Down. For example, as shown in the $(6,3)$-broom in Figure 3, $(u_{1},v_{1})$ is with the smallest partial order. Its direction is Down since it shares the same endpoint $v_{1}$ with the next edge $(u_{2},v_{1})$. Similarly, $(u_{2},v_{1})$, $(u_{3},v_{2})$, $(u_{4},v_{2})$, $(u_{5},v_{3})$ are also with direction Down, highlighted with blue arrows. $(u_{3},v_{1})$, $(u_{5},v_{2})$ are with direction Up, highlighted with reversed red arrows. $(u_{6},v_{3})$ has no direction since it is the edge with the maximum tuple.
@@
To begin with, we sort $U(G)$, $V(G)$ by the increasing order of colors $c(u),c(v)$ (Algorithm 2), and sort $E(G)$ by the increasing order of the color tuple $(c(u),c(v))$ (Algorithm 2).
Let $Dp[len][pre]$ denote the number of possible ways to build a growing broom with the current number of edges lenlen and the last edge prepre, which follows the fixed structure and order of edges by B. The objective is to correctly computed all states for $1\leq len\leq p+q-1$ (there are p+q1p+q-1 edges in total besides the first one) and $pre\in E(G)$. We use a 2D table to store the results, initialized to 0 (Algorithm 2). The base case contains the case where each $(u,v)$ acts as the first edge in the broom: we set $Dp[1][(u,v)]$ to be $1$ (Algorithm 2, 2). We iterate this dimension from small to large to accumulate the growing broom from $1..t$-th edges to $1..t+1$-th edges (Algorithm 2).
We use $Dir[t]$ to denote the $t$-th edge’s direction in the $(p,q)$-broom. When $Dir[t]$ is Up (Algorithm 2), the $t+1$-th edge and the $t$-th edge share a common point in $U$. Therefore, for each edge $(u,v)$ (Algorithm 2), $Dp[t+1][(u,v)]$ should accumulate the number of ways coming from any $Dp[t][(u,w)]$ such that $w$ is in $N(u,G)$ and $c(w)<c(v)$ (Algorithm 2). Similarly, when $Dir[t]$ is Down (Algorithm 2), the $t+1$-th edge and the $t$-th edge share a common point in $V$. In this case, $Dp[t+1][(u,v)]$ should accumulate the number of ways coming from any $Dp[t][(w,v)]$ such that $w$ is in $N(v,G)$ and $c(w)<c(u)$ (Algorithm 2, 2). After all computations, we return $B$ as the sum of all $Dp[p+q-1][\cdot]$ and the $Dp$ table as the results (Algorithm 2).
@@
Output: DpDp: a $2D$ array; $B$: the total number of $(p,q)$-brooms in $G$.
@@ 3 Sort $E(G)$ by the increasing order of $(c(u),c(v))$;
5Initialize DpDp with zeros;
@@ 4.3. Approximate Counting via Sampling
After CountingIndex(G,p,q)(G,p,q), we have accumulated $B$ $(p,q)$-brooms. To approximate the quantity relation between our computed $(p,q)$-brooms and $(p,q)$-bicliques, we can do the following sampling:
@@
After applying this sampling process sufficiently many times, we can roughly approximate the number of $(p,q)$-bicliques by $cnt_{biclique}/cnt_{broom}\times B.$ However, such a sampling process is too inefficient. We now propose Algorithm 3 to accelerate this process, utilizing the computed dynamic programming results from CountingIndex(G,p,q)(G,p,q). Recall that $Dp[t][(u,v)]$ represents the number of ways to build a growing $(p,q)$-broom til the $t$-th edge, which is $(u,v)$. To begin with, we start by sampling the last edge of the $(p,q)$-broom, denoted as $(u_{last},v_{last})$ (Algorithm 3). Here, the sampling should follow the weight distribution of $\{Dp[p+q-1][\cdot]\}$, indicating how many $(p,q)$-brooms end with each edge. The following process is similar to reverting the dynamic programming process. Instead of building the $(p,q)$-broom from the $1$-st edge to the (p+q1)(p+q-1)-th edge, we do it reversely. We use $(u_{cur},v_{cur})$, $U^{\prime}$, $V^{\prime}$ to denote the current growing $(p,q)$-broom. They are initialized as $(u_{last},v_{last})$, $\{u_{last}\}$, $\{v_{last}\}$ respectively (Algorithm 3, 3). Specifically, since $(u_{last},v_{last})$ is the (p+q1)(p+q-1)-th edge, we are now adding the (p+q2)(p+q-2)-th edge til the $1$-th edge gradually (Algorithm 3).
Assume we are processing the $i$-th edge now, and we know the $(i+1)$-th edge is $(u_{cur},v_{cur})$. We initialize an edge set $S$ to store the candidate edge for the $i$-th edge. Based on the direction signaling, we can know whether it should share the common endpoint in $U$ or $V$. If $Dir[i]$ is Up (Algorithm 3), then we should find all $w\in N(u_{cur},G)$ such that $c(w)<c(v_{cur})$ and $U^{\prime}\subseteq N(w,G)$. We assign $S$ to be $\{(u_{cur},w)\}$ (Algorithm 3). The first condition is for the color order, and the second condition indicates that adding $w$ can guarantee that the induced subgraph is biclique. Similarly, if $Dir[i]$ is Down (Algorithm 3), we find all all $w\in N(v_{cur},G)$ such that $c(w)<c(u_{cur})$ and $V^{\prime}\subseteq N(w,G)$. Then we assign $S$ to be $\{(w,v_{cur})\}$ (Algorithm 3).
@@
The idea to accelerate the sampling is, by building the $(p,q)$-broom, we strictly guarantee that it will correspond to a $(p,q)$-biclique while keeping track of the probability of sampling out this $(p,q)$-broom, which can be computed through the dynamic programming table. In the beginning, we initialize ansans to be $1$ (Algorithm 3). Then, for the $i$-th edge, the probability contribution of sampling it out from $S$ should be the number of the growing broom ends with $(u,v)\in S$ divided by the total number of possible brooms at this step, which is $\frac{\sum_{(u,v)\in S}Dp[i][(u,v)]}{Dp[i+1][(u_{cur},v_{cur})]}.$ We multiply ansans by this value (Algorithm 3). If ansans becomes 0, then there is no possible $(p,q)$-biclique from the current sampled $(p,q)$-broom. We return with 0 in this case (Algorithm 3). Now we are ready to sample the $i$-th edge from $S$ (Algorithm 3). We do so following the normalized distribution of possible brooms from this step ($\{Dp[i][\cdot]\}$). To continue the process, we assign $(u_{cur},v_{cur})$, $U^{\prime}$, $V^{\prime}$ to $(u_{next},v_{next})$, $U^{\prime}\cup\{u_{next}\}$, $V^{\prime}\cup\{v_{next}\}$, respectively (Algorithm 3, 3).
@@
In the end, we successfully sample a $(p,q)$-broom that corresponds to one $(p,q)$-biclique and compute the probability of sampling it out. We return this probability by multiplying $B$ as the approximate count of $(p,q)$-bicliques (Algorithm 3). It is not hard to see that the time complexity is dominated by the cost of updating $S$, which requires enumerating all neighbors. The loop only lasts $O(p+q)$. Therefore, the time complexity is $O((p+q)\times\Delta)$, where $\Delta$ denotes the maximal degree in $G$.
@@
Input: $G$: a colored bipartite graph; $p,q$: two parameters; DpDp: a $2D$ array; $B$: the total number of $(p,q)$-brooms in $G$.
Output: ansans: estimate number of $(p,q)$-clique in $G$.
@@ 19  $U^{\prime}\leftarrow U^{\prime}\cup\{u_{next}\},V^{\prime}\leftarrow V^{\prime}\cup\{v_{next}\}$;
4.4. Overall Algorithm
We provide Algorithm 4 as a black box to call all three subroutines properly and output the estimated answer of the biclique count. We start by coloring (Algorithm 4) and pattern counting (Algorithm 4). Then by the input sampling size parameter $T$, we repeatedly call Sampling(G,p,q,Dp,B)(G,p,q,Dp,B) and aggregate the return to $\hat{C}$. The output approximate $(p,q)$-biclique count is the average of all $T$ attempts (Algorithm 4). Note that as a common trick, we will execute core-reduction for the original graph (Yang et al., 2021): For any query $(p,q)$, we split the graph and reduce the query pair to (p1,q)(p-1,q). Now, we give a theoretical analysis of CBS in the following.
@@ 4.4. Overall Algorithm
1 $G\leftarrow\$Coloring(G,p,q)(G,p,q);
2 $Dp,B\leftarrow\$CountingIndex(G,p,q)(G,p,q);
@@ 5  $\hat{C}\leftarrow\hat{C}\$+ Sampling(G,p,q,Dp,B)(G,p,q,Dp,B);
4.4.1. Unbiasedness.
We now show that the CBS method is unbiased. Let $(u_{cur},v_{cur})$, $U^{\prime}$, and $V^{\prime}$ represent the current growing $(p,q)$-broom. For clarity, we introduce the following definitions:
  • $\mathcal{F}_{U^{\prime},V^{\prime}}$: The total value to be multiplied into ansans, defined as the product of the last $i=(p+q)-(|U^{\prime}|+|V^{\prime}|)$ fractions.
@@
(a) When $|U^{\prime}|+|V^{\prime}|=p+q$, we have $\mathcal{B}_{U^{\prime},V^{\prime}}=1$. In this case, $\mathcal{F}_{U^{\prime},V^{\prime}}=\mathcal{C}_{U^{\prime},V^{\prime}}=1$ if $U^{\prime},V^{\prime}$ forms a $(p,q)$-biclique; otherwise $\mathcal{F}_{U^{\prime},V^{\prime}}=\mathcal{C}_{U^{\prime},V^{\prime}}=0$. Hence, $\mathbb{E}\left[\mathcal{F}_{U^{\prime},V^{\prime}}\right]=\mathcal{C}_{U^{\prime},V^{\prime}}/\mathcal{B}_{U^{\prime},V^{\prime}}$ trivially holds in this case.
(b) Suppose that the lemma holds for all growing $(p,q)$-brooms with $|U^{\prime}|+|V^{\prime}|=k+1$ . For any broom with $|U^{\prime}|+|V^{\prime}|=k$, let $P=\frac{\sum_{(u,v)\in S}Dp[i][(u,v)]}{Dp[i+1][(u_{cur},v_{cur})]}.$ We then have
@@
Since $u_{cur}=u$ or $v_{cur}=v$ always holds, we have that $|U^{\prime}\cup\{u\}|$$|V^{\prime}\cup\{v\}|$=|U|+|V|+1=k+1.=|U|+|V|+1=k+1. According to the definition of $Dp$ and $\mathcal{B}$, We can derive that $Dp[i][(u,v)]=\mathcal{B}_{U^{\prime}\cup\{u\},V^{\prime}\cup\{v\}}$, and thus
@@
$=\frac{\sum\limits_{(u,v)\in S}\mathcal{C}_{U^{\prime}\cup\{u\},V^{\prime}\cup\{v\}}}{\mathcal{B}_{U^{\prime},V^{\prime}}}=\frac{\mathcal{C}_{U^{\prime},V^{\prime}}}{\mathcal{B}_{U^{\prime},V^{\prime}}},$
@@
Let $ans_{i}$ be the value of ansans in the $i$-th sampling. Let $\hat{C}=\frac{1}{T}\sum_{i=1}^{T}ans_{i}\times B$. Then $\hat{C}$ is an unbiased estimator of the number of $(p,q)$-bicliques in $G$.
@@
Then, we can derive that: $\mathbb{E}\left[\hat{C}\right]=\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}\left[ans_{i}\times B\right]=\mathcal{C}_{\emptyset,\emptyset}.$
Note that $\mathcal{C}_{\emptyset,\emptyset}$ is the number of $(p,q)$-bicliques in $G$ according to the definition. Therefore, $\hat{C}$ is an unbiased estimator of the number of $(p,q)$-bicliques in $G$. ∎
@@
4.4.2. Error Analysis.
we now analyze the estimation error of our sampling algorithms. Our analysis relies on the classic Hoeffding’s inequalities, which are shown below.
@@
For the random variables $X_{i}\in[0,M]$$1\leq i\leq n$, we let $X=\sum_{i=1}^{n}X_{i}$. Then for $\epsilon>0$, we have
@@
We can derive $ans_{i}\times B\in[0,B]$ since the values multiplied to ansans are all probabilities in $[0,1]$. And we have $\mathbb{E}\left[\hat{C}T\right]=CT$ based on Theorem 4.4. Given a positive value $\epsilon$, applying Lemma 4.5 by plugging in $n=T$, $X_{i}=ans_{i}\times B$, $X=\sum_{i=1}^{n}X_{i}=\hat{C}T$, we then have
@@
Further, we have $\Pr(\frac{|\hat{C}-C|}{C}\geq\epsilon)\leq 2\exp(-\frac{2\epsilon^{2}C^{2}T}{B^{2}}).$
Let $2\exp(-\frac{2\epsilon^{2}C^{2}T}{B^{2}})\leq\alpha,$ we can derive that $T\geq\frac{B^{2}}{2\epsilon^{2}C^{2}}\ln(\frac{2}{\alpha}).$
4.4.3. Time and Space Complexity.
The time complexity for each of the three parts in CBS are already discussed separately. We now give the following theorem as a summary.
Theorem 4.7.
The time complexity of CBS is $O(\kappa|E|+(p+q)|E|+T\cdot(p+q)\cdot\Delta)$, where $\Delta$ denotes the maximal degree in $G$. Specifically, to reach a $(1+\epsilon)$-approximation with probability $(1-\alpha)$, we require $T\geq\frac{B^{2}}{2\epsilon^{2}C^{2}}\ln(\frac{2}{\alpha})$ where $C,B$ denotes the number of $(p,q)$-bicliques and $(p,q)$-brooms.
Proof.
The first two terms follow from the analysis in Section 4.1 and Section 4.2. The third term follows from the fact that we need to conduct the sampling process (Algorithm 3) for $T$ times, and each time costs $(p+q)\cdot\Delta$. The approximation guarantee comes directly from Theorem 4.6. ∎
We now provide the memory usage analysis as follows. We have also conducted an empirical experiment on the memory consumption in Section 6.4.
Theorem 4.8.
The space complexity of CBS is $O(|U|+|V|+(p+q)\cdot|E|)$.
Proof.
In Algorithm 1, since we need to maintain $Cnt$ and $S_{next}$ which is at most $|U|+|V|$, the memory usage is $O(|U|+|V|)$. In Algorithm 2, the memory usage is dominated by the size of the DP table, thus $O((p+q)\cdot|E|)$. Algorithm 3 is for sampling a $(p,q)$-biclique, and $|S|$ is at most $\max\deg(u)$, thus the memory usage is $O(\max\deg(u))$. In all, the total space complexity is $O(|U|+|V|+(p+q)\cdot|E|)$. ∎
The EP/Zz++ uses the heuristic framework to combine EPivoter (for exact counting on small portion of graphs) and ZigZag++. In the analysis of ZigZag++, its time and space complexity is $O((p+q)|E|+(p+q)^{2}|V|+T\cdot\Delta\cdot(p+q)^{2})$ and $O((p+q)|E|)$. Our memory usage is similar to each other. If we consider $T$ and $\kappa$ as a constant, our time complexity has some advantage as we don’t have the term of $(p+q)^{2}$. However, it is obvious that both our algorithms’ performance are highly determined by the size of $T$.
From Theorem 4.6, the sample size $T$ is mainly determined by $(\frac{B}{C})^{2}$. That is to say, we need a larger sample size $T$ to ensure high accuracy with a larger $(\frac{B}{C})^{2}$, where $B$ is the number of $(p,q)$-brooms, $C$ is the number of $(p,q)$-bicliques. As shown in Section 6.3, $(\frac{B}{C})^{2}$ is usually small comparing to the dominating term of EP/Zz++ ($\frac{Z^{2}}{\rho^{2}}$(Ye et al., 2023). Therefore, our algorithm generally does not need a large $T$ to achieve good accuracy in real-world datasets. It can also be concluded that CBS guarantees an equivalent error ratio with significantly fewer sampling times and higher precision with the same sampling times compared to EP/Zz++.